Current stock price is $1,000. The market expects the price will be $1,700 in 2 years. What should be the return over the next two years if return is continuously compounded?
Answer #1

FV = PV*e^(r*t)

1700 = 1000*e^(2*r)

ln(1700/1000)=2*r

0.5306=2*r

r=26.53%